can someone please help me with this question ? A water wheel rotates through the angle theta,the water level L behind the wheel changes according to the equation: L =2-2sintheta-4cos^2theta where l is measured in inches. determine the values of theta for which the water level is zero. find the exact values. do not use a claculator
Answer by Alan3354(69443)

L =2 - 2sin(t) - 4cos^2(t) = 0
1 - sin(t) - 2cos^2(t) = 0
1 - sin(t) - 2(1 - sin^2(t)) = 0
2sin^2 - sin - 1 = 0
(2sin + 1)*(sin - 1) = 0
sin = 1 --> theta = 90 + n*180 degs, n = 0,1,2,3...
----
sin = -1/2 --> theta = 210 + n*360 degs, n = 0,1,2,3...
sin = -1/2 --> theta = 330 + n*360 degs, n = 0,1,2,3...